# DynoSAM: Open-Source Smoothing and Mapping Framework for Dynamic SLAM

### v2

@@ V Foundations for Dynamic SLAM
Under the standard assumption of zero-mean Gaussian noise, we take the negative log of Eq. (11) and drop the scaling-factor allowing us to collect all residuals and construct a nonlinear least-squares problem:
$\mathbf{\theta}^{\text{MAP}}=\underset{\mathcal{X}}{\mathrm{argmin}}\sum\lVert{\mathbf{r}}\rVert^{2}_{\Sigma}$ (12)
For completeness, our framework also includes the object-centric formulations explored in [18]. However, due to their inferior performance, this work does not discuss them further.
V-A Measurements
We define the following notation for visual features. Direct measurements are denoted as $\mathbf{z}_{\text{2D}}\in\mathbb{R}^{2}$ and represent 2D pixel measurements on the image plane. We denote 3D point measurements in the local sensor frame $\{X\}$ as $\mathbf{z}_{\text{3D}}\in\mathbb{R}^{3}$. In the context of stereo/RGB-D SLAM, we assume the depth $d$ of each pixel measurement $\mathbf{z}_{\text{2D}}$ is available for every time-step $k$, and 3D measurements are constructed accordingly:
@@ V-A Measurements
where $\pi^{-1}(\cdot)$ is the sensors back-projection function, e.g. pinhole camera model. We further specify the notation for sets of measurements corresponding to static and dynamic entities. $\mathcal{S}_{\text{2D},k}$ and $\mathcal{S}_{\text{3D},k}$ denote the set of static 2D keypoints $\mathbf{z}_{\text{2D}}$ and their corresponding 3D landmark measurements $\mathbf{z}_{\text{3D}}$ at time $k$, respectively. Similarly, $\mathcal{D}_{\text{2D},k}$ and $\mathcal{D}_{\text{3D},k}$ are the sets of dynamic measurements. The set of all static measurements (2D and 3D) is denoted as $\mathcal{S}_{k}$. $\mathcal{D}_{k}$ is the set of all dynamic measurements. The set of dynamic measurements on a particular object is specified using the super-script $j$, e.g. $\mathcal{D}_{\text{3D},k}^{j}$ is the sub-set of 3D measurements on object $j$ at time $k$.
V-B World-Centric Motion Estimator
The world-centric motion formulation directly estimates for object motions and their rigid-structure, as well as the static scene and camera pose:
@@ V-B World-Centric Motion Estimator
This formulation was first proposed in [15] and then rigorously evaluated in [18], establishing it as a highly accurate method for object motion estimation. Fig. 3 shows the corresponding factor graph for this formulation and includes three static landmarks (blue) and one landmark (orange) on a dynamic object that are tracked over three frames.
@@
where the odometry $\prescript{X_{k-1}}{k-1}{\mathbf{T}}_{k}$ can be estimated by the front-end of a visual SLAM system, such as the proposed one described in Sec. VI-B, or from alternative sources such as an Inertial Measurement Unit (IMU) if available. The operation $\left[\log\left(\cdot\right)\right]^{\vee}$ maps an $\mathrm{SE}(3)$ transformation to $\mathbb{R}^{6}$.
From Eq. (10) we derive the ternary object motion factor which models the motion of any point $i$ on a rigid body $j$ as in [15, 18]:
@@
This enforces a constant motion model represented in $\{W\}$ and prevents abrupt, drastic and unrealistic changes in object motions between consecutive frames. Previous work [11] shows that if the motion is constant in the body frame, then the reference frame pose change is also constant; an extended version of this proof is included in Appendix -B.
@@
$+$ $\sum_{m\in\mathcal{S}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}$
$+$ $\sum_{m\in\mathcal{D}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}$
$+$ $\sum_{j\in\mathcal{J}_{k}}\>\sum_{m\in\mathcal{D}_{\text{3D},k}^{j}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},H}}\rVert^{2}_{\Sigma_{{{\text{3D},H}}}}$
$+$ $\sum_{j\in\mathcal{J}_{k}}\lVert{\mathbf{r}_{\delta H}}\rVert^{2}_{\Sigma_{{{\delta H}}}}\biggr{)}\text{,}$
where $\mathbf{r}_{0}$ is prior residual on the first camera pose state variable and $\rho_{h}$ is a robust cost function. In our implementation we use the Huber [53] function, although any robust cost can be used.
While not being directly incorporated in this estimation, the trajectory of each object ${{}^{W}\mathcal{L}}$ can be recovered by recursively propagating the previous pose using the estimated motion following Eq. (9). This method requires an arbitrary first pose $\prescript{W}{}{\mathbf{L}}_{s}$ to be defined, which anchors the resulting trajectory. Since the estimated motion is an absolute motion, we can define this pose anywhere relative to $\{W\}$. Practically, we construct $\prescript{W}{}{\mathbf{L}}_{s}$ upon the initial observation of an object using the centroid of all object point observations as its position and identity as its orientation. However, any method that provides object pose, such as learned techniques can easily be adopted. For evaluation, we use the ground truth object pose to set $\prescript{W}{}{\mathbf{L}}_{s}$. This ensures that the estimated object trajectories and the ground truth have aligned origins which facilitates proper evaluation as discussed in Sec. VII-A.
V-C World-Centric Pose Estimator
While estimating $\prescript{W}{k-1}{\mathbf{H}}_{k}$ provides valuable information about object dynamics, directly estimating object pose $\prescript{W}{}{\mathbf{L}}_{k}$ can offer a more comprehensive understanding of the scene. As seen in the literature, object pose enables more direct reasoning about the semantic and geometric properties of objects, facilitating tasks such as scene understanding and object-centric planning. We therefore further propose a novel approach that directly estimates for the object pose:
@@ V-C World-Centric Pose Estimator
The core of this approach is as follows: Eq. (8) shows how a motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$ explicitly models a pair of consecutive object poses. Using this idea, we reformulate our factors to make every $\prescript{W}{}{\mathbf{L}}$ a variable of the system. The corresponding factor graph is shown in Fig. 4. Substituting Eq. (8) into Eq. (17) forms a new quaternary object motion factor:
@@
We similarly define the smoothing factor in-terms of object pose by substituting Eq. (8) into Eq. (18). This forms the object pose smoothing factor:
@@
$\begin{split}\mathbf{\theta}^{\text{MAP}}=&\underset{\bm{\mathbf{\theta}}}{\mathrm{argmin}}\biggl{(}\lVert{\mathbf{r}_{0}}\rVert^{2}_{\Sigma_{{0}}}+\lVert{\mathbf{r}_{\delta X}}\rVert^{2}_{\Sigma_{{\delta X}}}\\ +&\sum_{s\in\mathcal{S}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}\\ +&\sum_{d\in\mathcal{D}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}\\ +&\sum_{j\in\mathcal{J}_{k}}\>\sum_{m\in\mathcal{D}_{\text{3D},k}^{j}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},LL}}\rVert^{2}_{\Sigma_{{{\text{3D},LL}}}}\\ +&\sum_{j\in\mathcal{J}_{k}}\lVert{\mathbf{r}_{\delta LL}}\rVert^{2}_{\Sigma_{{{\delta LL}}}}\biggr{)}\end{split}$ (23)
@@
Unlike other Dynamic SLAM methods which also estimate for object pose, our formulation continues to model dynamic points directly in the world frame rather than using an object-centric representation. This avoids the problems of object-centric formulations, as outlined in [18]. Furthermore, by simply re-parameterizing our cost functions in terms of pose we can ensure that our formulation directly embeds the observed motion model, which allows us to define $\{L\}$ anywhere on the object. This flexibility enables the integration of various initialization methods, such as SLAM front-ends or observation centroids, without compromising rigid-body kinematic constraints.
VI System
@@
Figure 5: DynoSAM system diagram. Our pipeline receives RGB, Segmentation, Optical Flow and Dense Depth images (on the left) as input and outputs static/dynamic map points, object and robot trajectories. A complete front-end is included for feature and object tracking as well as a back-end where different Dynamic SLAM formulations are implemented.
Based on the formulations discussed in Sec. V, we present the DynoSAM pipeline, a factor-graph based Dynamic Visual SLAM system. For input, our pipeline takes stereo/RGB-D, instance segmentation masks and dense optical flow, and produces globally consistent trajectories of the camera and objects as well as the static structure and temporal map of each object. The system diagram is shown in Fig. 5.
@@
Our pipeline is broken into the typical frontend-backend structure. The front-end (Sec. VI-B) processes image data to generate static and dynamic feature tracks and provides initial estimates for the back-end. The front-end is responsible for object-level data association across frames, ensuring that features are consistently tracked on the same object to enable robust estimation. The back-end (Sec. VI-C) fuses static and dynamic measurements via a GTSAM-based [54] factor-graph optimization to produce a globally consistent estimate of the dynamic scene. Our implementation can perform full-batch smoothing where the system is optimized over all measurements, or a sliding-window estimation which bounds the size of the optimization problem.
@@
We implement both the world-centric motion and pose formulations explained in Sec. V as part of our system, and exhaustively test and analyze them on a wide variety of datasets, the result of which is presented in Sec. VII.
VI-A Image Preprocessing
@@
The front-end requires a set of four images $\mathcal{I}_{k}$ as input per time-step. Each $\mathcal{I}_{k}$ consists of a RGB image $I^{R}_{k}$, an instance segmentation $I^{S}_{k}$, per-pixel optical-flow $I^{F}_{k}$, and per-pixel depth $I^{D}_{k}$, as shown in Fig. 5. We undistort and align $I^{R}_{k}$with the depth image $I^{D}_{k}$that is obtained from an RGB-D or stereo camera system. Both depth and optical-flow are expected to be dense and can be obtained from classical [55] or learned methods [56]. Finally, $I^{S}_{k}$is used in the process of masking out dynamic objects from the static background. Our system requires all background pixels to be labeled as 00 and object pixels in the scene are labeled ${1\dots n_{o}}$. This processing is performed online as an integral part of our complete SLAM pipeline, using built-in algorithms to compute dense optical flow and semantic instance segmentation from $I^{R}_{k}$.
VI-B Front-end
@@
The DynoSAM front-end performs per-frame visual tracking, producing an initial estimation of camera pose $\prescript{W}{}{\mathbf{X}}_{k}$ and per-object motions $\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k}$ as well as a set of visual measurements as defined in Sec. V-A:
@@
The reminder of this section introduces each front-end module. The Feature Detection & Tracking module (Sec. VI-B1) generates $\mathcal{Z}_{k}$ by tracking static and dynamic features from $\mathcal{I}_{k}$. These measurements support camera pose estimation (Sec. VI-B2) and multi-object motion estimation (Sec. VI-B3). Joint Optical-Flow (Sec. VI-B4) and Object Motion Refinement (Sec. VI-B5) components are used to further refine camera pose and object motion estimates.
VI-B1 Feature Detection & Tracking
The feature detection and tracking module matches corresponding static and per-object features between consecutive frames. $I^{R}_{k}$and $I^{F}_{k}$are used for feature detection and tracking while $I^{S}_{k}$is used to determine if a keypoint belongs to the static background or a dynamic object. The module ensures consistent tracking of features on the same object while maintaining sufficient spread and density of features across frames, thereby preserving the accuracy of both visual odometry and object motion estimation.

### v3

@@ IV Foundations for Dynamic SLAM
Under the standard assumption of zero-mean Gaussian noise, we take the negative log of (10) and drop the scaling-factor allowing us to collect all residuals and construct a nonlinear least-squares problem:
$\mathbf{\theta}^{\text{MAP}}=\underset{\mathcal{\theta}}{\mathrm{argmin}}\sum\lVert{\mathbf{r}}\rVert^{2}_{\Sigma}$ (11)
For completeness, our framework also includes the object-centric formulations explored in [morris2024icra]. However, due to their inferior performance, this work does not discuss them further.
IV-A Measurements
We define the following notation for visual features. Direct measurements are denoted as $\mathbf{z}_{\text{2D}}\in\mathbb{R}^{2}$ and represent 2D pixel measurements on the image plane. We denote 3D point measurements in the local sensor frame $\{X\}$ as $\mathbf{z}_{\text{3D}}\in\mathbb{R}^{3}$. In the context of stereo/RGBD SLAM, we assume the depth $d$ of each pixel measurement $\mathbf{z}_{\text{2D}}$ is available for every time-step $k$, and 3D measurements are constructed accordingly:
@@ IV-A Measurements
where $\pi^{-1}(\cdot)$ is a sensor’s back-projection function, e.g. pinhole camera model. We further specify the notation for sets of measurements corresponding to static and dynamic entities. $\mathcal{S}_{\text{2D},k}$ and $\mathcal{S}_{\text{3D},k}$ denote the set of static 2D keypoints $\mathbf{z}_{\text{2D}}$ and their corresponding 3D landmark measurements $\mathbf{z}_{\text{3D}}$ at time $k$, respectively. Similarly, $\mathcal{D}_{\text{2D},k}$ and $\mathcal{D}_{\text{3D},k}$ are the sets of dynamic measurements. The set of all static measurements (2D and 3D) is $\mathcal{S}_{k}$. The set of dynamic measurements on a particular object is specified using the super-script $j$, e.g. $\mathcal{D}_{\text{3D},k}^{j}$ is the sub-set of 3D measurements on object $j$ at time $k$.
IV-B World-Centric Motion Estimator
The formulation for the World-Centric Motion Estimator (WCME) directly estimates object motions and their rigid-structure, as well as the static scene and camera pose:
@@ IV-B World-Centric Motion Estimator
This formulation was first proposed in [zhang2020vdoslam] and evaluated in [morris2024icra] which established it as a highly accurate method for object motion estimation. Fig. 3 shows the corresponding factor graph for this formulation and includes three static landmarks (blue) and one landmark (orange) on a dynamic object that are tracked over three frames.
@@
where the odometry $\prescript{X_{k-1}}{k-1}{\mathbf{T}}_{k}$ can be estimated by the front-end of a visual SLAM system, such as the proposed one described in Sec. V-B, or from alternative sources such as an Inertial Measurement Unit (IMU) if available. The operation $\left[\log\left(\cdot\right)\right]^{\vee}$ maps an $\mathrm{SE}(3)$ transformation to $\mathbb{R}^{6}$.
From (9) the ternary object motion factor models the motion of any point $i$ on a rigid body $j$ as in [zhang2020vdoslam, morris2024icra]:
@@
This enforces a constant motion model represented in $\{W\}$ and prevents abrupt, drastic and unrealistic changes in object motions between consecutive frames. This constraint holds since any constant change in the body frame also leads to a constant change in the reference frame pose change, as shown in [Henein20icra]; an extended version of this proof is included in Appendix -B.
@@
$+\sum_{m\in\mathcal{S}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}+\sum_{m\in\mathcal{D}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}$
$+\sum_{j\in\mathcal{J}_{k}}\>\sum_{m\in\mathcal{D}_{\text{3D},k}^{j}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},H}}\rVert^{2}_{\Sigma_{{{\text{3D},H}}}}+\sum_{j\in\mathcal{J}_{k}}\lVert{\mathbf{r}_{\delta H}}\rVert^{2}_{\Sigma_{{{\delta H}}}}\biggr)\text{,}$
where $\mathbf{r}_{0}$ is prior residual on the first camera pose state variable and $\rho_{h}$ is any robust cost function. In our implementation, we use the Huber [Huber92bs] function. While not being directly incorporated in this estimation, the trajectory of each object $^{W}\mathcal{L}$} can be recovered by recursively propagating the previous pose using the estimated motion following (8):
$\prescript{W}{}{\mathbf{L}}_{k}=\prescript{W}{k-1}{\mathbf{H}}_{k}\dots\prescript{W}{s}{\mathbf{H}}_{s+1}\>\prescript{W}{}{\mathbf{L}}_{s}\text{,}$ (19)
where $s$ is the first time-step at which this object is observed. This method requires an arbitrary first pose $\prescript{W}{}{\mathbf{L}}_{s}$ to be defined, which anchors the resulting trajectory. Since the estimated motion is an absolute motion, we can define this pose anywhere relative to $\{W\}$. Practically, we construct $\prescript{W}{}{\mathbf{L}}_{s}$ upon the initial observation of an object using the centroid of all object point observations as its position and identity as its orientation. In addition, any method that provides object pose, such as learned techniques, can easily be adopted. For experiment evaluation, we use the ground truth object pose to set $\prescript{W}{}{\mathbf{L}}_{s}$, aligning the origins of estimated object trajectories and that of the ground truth to facilitates proper evaluation as discussed in Sec. VI-B.
IV-C World-Centric Pose Estimator
Up to this point, we have used the rigid-body motion $\prescript{W}{}{\mathbf{H}}$ to directly parametrize the object state. This perspective, used in both our work and prior studies, has enabled effective estimation of object motion and recovery of object pose. However, while $\prescript{W}{k-1}{\mathbf{H}}_{k}$ provides valuable insight regarding object dynamics, directly estimating the object pose $\prescript{W}{}{\mathbf{L}}_{k}$ is highly relevant for downstream tasks. We therefore propose the World-Centric Pose Estimator (WCPE), a novel formulation that parametrizes the object state in terms of its pose $\prescript{W}{}{\mathbf{L}}$. Unlike other estimators that use a motion prior [judd2024ijrr_mvo, bescos2021ral] between poses, our formulation uses the rigid-body motion directly as a constraint between poses, based on the measurement of object points. This approach therefore demonstrates the importance of using the absolute motion representation as the foundation for new Dynamic SLAM formulations.
WCPE estimates the states:
@@ IV-C World-Centric Pose Estimator
The core of this approach is as follows: equation (7) shows how a motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$ explicitly models a pair of consecutive object poses. Using this idea, we reformulate our factors to make every $\prescript{W}{}{\mathbf{L}}$ a variable of the system. The corresponding factor graph is shown in Fig. 4. Substituting (7) into (16) forms a new quaternary object motion factor:
@@
We similarly define the smoothing factor in-terms of object pose by substituting (7) into (17). This forms the object pose smoothing factor:
@@
$\begin{split}\mathbf{\theta}^{\text{MAP}}=\underset{\bm{\mathbf{\theta}}}{\mathrm{argmin}}\biggl(\lVert{\mathbf{r}_{0}}\rVert^{2}_{\Sigma_{{0}}}+\lVert{\mathbf{r}_{\delta X}}\rVert^{2}_{\Sigma_{{\delta X}}}\\ +\sum_{s\in\mathcal{S}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}+\sum_{d\in\mathcal{D}_{\text{3D},k}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},R}}\rVert^{2}_{\Sigma_{{{\text{3D},R}}}}\\ +\sum_{j\in\mathcal{J}_{k}}\>\sum_{m\in\mathcal{D}_{\text{3D},k}^{j}}\rho_{h}\lVert{\mathbf{r}_{\text{3D},LL}}\rVert^{2}_{\Sigma_{{{\text{3D},LL}}}}+\sum_{j\in\mathcal{J}_{k}}\lVert{\mathbf{r}_{\delta LL}}\rVert^{2}_{\Sigma_{{{\delta LL}}}}\biggr)\end{split}$ (23)
@@
We highlight that our formulation represents both object motions and their points in the world frame. This differs from existing methods that also estimate for object pose but use an object-centric representation for points [bescos2021ral, morris2024icra]. This small difference is vital to ensure the rigid-body motion model remains explicit and therefore the potential problems of object-centric formulations outlined by Morris et al[morris2024icra] are avoided. Furthermore, by continuing to base our formulation on the absolute motion representation, we can define $\{L\}$ anywhere on the object, as the absolute motion representation is independent of the object’s reference frame and pose (Sec. III-A). This flexibility enables the integration of various initialization methods, e.g. SLAM front-ends or observation centroids, without compromising rigid-body kinematic constraints.
V System
@@
Figure 5: DynoSAM system diagram. Our pipeline receives RGB, Segmentation, Optical Flow and Dense Depth images (on the left) as input and outputs static/dynamic map points, object and robot trajectories, shown in (a). A complete front-end is included for feature and object tracking as well as a back-end where different Dynamic SLAM formulations are implemented. The per-object map and trajectory is highlighted in (b) and the camera trajectory in (c).
Our work is primarily focused on the estimation of object motions and poses and it is highly relevant to provide an open-source Dynamic SLAM system to facilitate evaluation and wide-spread use. Based on the formulations discussed in Sec. IV, this section presents the DynoSAM pipeline, a factor-graph-based Dynamic Visual SLAM system.
@@
Our pipeline is broken into the typical frontend-backend structure. As shown in Fig. 5, our pipeline takes stereo/RGBD, instance segmentation masks and dense optical flow as input, and produces globally consistent trajectories of the camera and objects as well as the static structure and temporal map of each object (Fig. 5 (a-c)). The front-end (Sec. V-B) processes this input data to generate static and dynamic feature tracks and provides initial estimates for the back-end. It is responsible for object-level data association across frames, ensuring that features are consistently tracked on the same object to enable robust estimation. The robust handling of noisy depth and optical flow data, as well as segmentation inaccuracies, have been well studied in et al[Henein20icra, zhang20iros], which serve as inspiration for our front-end implementation. Further evaluations, while relevant for real-world deployment, is beyond the scope of this particular work, since our primary focus is the understanding and evaluation of the Dynamic SLAM back-end. The back-end (Sec. V-C) implements our presented formulations in Sec. IV and defines the Dynamic SLAM estimation problem. It fuses static and dynamic measurements via a GTSAM-based [gtsam] factor-graph optimization to produce a globally consistent estimate of the dynamic scene. Our implementation is flexible to accommodate any custom formulation for Dynamic SLAM and options for full-batch smoothing where the system is optimized over all measurements, as well as a sliding-window estimation which bounds the size of the optimization problem are available. We emphasize that while our primary contributions encompass the back-end of a Dynamic SLAM pipeline, we highlight that our open-source front-end also includes implementation improvements, such as multi-threaded object tracking for improved efficiency, and a range of user options.
@@
We implement both the world-centric motion and pose formulations explained in Sec. IV as part of our system, and exhaustively test and analyze them on a wide variety of datasets, the result of which is presented in Sec. VI.
V-A Image Preprocessing
@@
The front-end requires a set of four images $\mathcal{I}_{k}$ as input per time-step. Each $\mathcal{I}_{k}$ consists of a RGB image $I^{R}_{k}$ , an instance segmentation $I^{S}_{k}$ , per-pixel optical-flow $I^{F}_{k}$ , and per-pixel depth $I^{D}_{k}$ , as shown in Fig. 5. We undistort and align $I^{R}_{k}$ with the depth image $I^{D}_{k}$ that is obtained from an RGBD or stereo camera system. Both depth and optical-flow are expected to be dense and can be obtained from classical [farneback2003denseflow] or learned methods [teed2020raft, huang2022flowformer]. Finally, $I^{S}_{k}$ is used in the process of masking out dynamic objects from the static background. Our system requires all background pixels to be labeled as 0 and object pixels in the scene are labeled $1\dots n_{o}$}. This processing is performed online as an integral part of our complete SLAM pipeline; currently RAFT [teed2020raft] and YOLOv8 [JocherUltralyticsYOLO2023] are used to compute dense optical-flow and semantic instance segmentation respectively.
V-B Front-end
@@
The DynoSAM front-end performs per-frame visual tracking, producing an initial estimation of camera pose $\prescript{W}{}{\mathbf{X}}_{k}$ and per-object motions $\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k}$ as well as a set of visual measurements as defined in Sec. IV-A:
@@
The reminder of this section introduces each front-end module. The Feature Detection & Tracking module (Sec. V-B1) generates $\mathcal{Z}_{k}$ by tracking static and dynamic features from $\mathcal{I}_{k}$. These measurements support camera pose estimation (Sec. V-B2) and multi-object motion estimation (Sec. V-B3). Joint Optical-Flow (Sec. V-B4) and Object Motion Refinement (Sec. V-B5) components are used to further refine camera pose and object motion estimates.
V-B1 Feature Detection & Tracking
The feature detection and tracking module matches corresponding static and per-object features between consecutive frames. $I^{R}_{k}$ and $I^{F}_{k}$ are used for feature detection and tracking while $I^{S}_{k}$ is used to determine whether a keypoint belongs to the static background or a dynamic object. This module ensures consistent tracking of features on the same object while maintaining sufficient spread and density of features across frames to preserve the accuracy of both visual odometry and object motion estimation.